OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x.f(x) = quantity x minus nine divided by quantity x plus five. and g(x) = quantity negative seven x minus eight divided by quantity x minus one.

OpenStudy (anonymous):

To clarify: \[\large \begin{cases}f(x)=\dfrac{x-9}{x+5}\\\\\\ g(x)=\dfrac{-7x-8}{x-1}\end{cases}\] Is that right?